When 12.0 mL of a 8.80E-4 M zinc acetate solution is combined with 25.0 mL of a 6.86E-4 M potassium carbonate solution does a precipitate form? (Y/N)

For these conditions the Reaction Quotient, Q, is equal to___.
Answer #1

ZafHxcee), + ķGo → Zalozts+ KC Hycoo (@ay Zn los d ant² + Cogna Ksp = [2n+27 [ coz ? 1 = 1-46X10-106ct 2981) Most nitrate andQ = _ - - 2n+27 [663 4.635x164 x 2.854x10-4 1323 x 10-7 a) ke Il because oc=1323x10 7 e ke=1.46x10 Qe=Ke Quy k oc<ke At equi